OpenStudy (anonymous):

One large pipe with cold water and one small pipe with hot water are used to fill a pool. When both pipes are used, the pool can be filled in 2 hours. If only the cold water pipe (large one) is used to fill the pool it can be done in 3 hours. How long would it take to fill the pool with just the hot water pipe?

OpenStudy (anonymous):

is this problem missing some more information ?

OpenStudy (anonymous):

that is all i got in the problem

OpenStudy (anonymous):

1/3 + 1/x = 1/2

OpenStudy (mertsj):

It takes the cold water pipe 3 hours to fill the pool. So it does 1/3 of the job per hour It takes the hot water pipe x hours to fill the pool. So it does 1/x of the job per hour. It takes both pipes 2 hours to fill the pipe so together they do 1/2 of the job per hour. 1/3+1/x=1/2 just like dockworker said.

OpenStudy (anonymous):

so it takes 6 hours to fill the pool witht hot water?

OpenStudy (anonymous):

???

OpenStudy (mertsj):

yes